OpenStudy (anonymous):

Solve using transformations. 24y = 96 answer: y = 4 is that correct? I have a couple of other questions too

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

Thanks, are you available for some more questions?

OpenStudy (anonymous):

a couple

OpenStudy (anonymous):

Great, the next one is 18 = 3a/4 the answer is a = 24 is that correct?

OpenStudy (anonymous):

That is correct!

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

Thanks! The next one is... 8(y - 10) + 2y = 20 y = 10 is that correct?

OpenStudy (anonymous):

Yes that is also correct

OpenStudy (anonymous):

Thank you, is it ok to keep going?

OpenStudy (anonymous):

Sure if you would like

OpenStudy (anonymous):

Thanks for the medal by the way, and thanks

OpenStudy (anonymous):

No problem, you are getting all your answers right, it seems that you understand it perfectly.

OpenStudy (anonymous):

I hope so, math isn't my best subject :P b - 4 = 4 + 7b/3 the answer is b = -6/8 is that correct?

OpenStudy (anonymous):

Hm give me a second

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

i did not get -6/8

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

That is strange I got -2

OpenStudy (anonymous):

b - 4 = 4 + 7b/3 3b - 12 = 12 + 7b -24 = 4b -6 = b

OpenStudy (anonymous):

Can you guys show me how you did it?

OpenStudy (anonymous):